(x-y)p-(x-y+z)q=z


1
Expert's answer
2021-03-31T16:53:35-0400

given,

"(x-y)p-(x-y+z)q=z"


This equation can be written as-

"(x-y)p+(y-x-z)q=z"


The lagrange's subsidary equation are-


"\dfrac{dx}{x-y}=\dfrac{dy}{y-x-z}=\dfrac{dz}{z}~~~~~~-(1)"

Each ratio of (1) is equal to-


"\dfrac{dx+dy+dz}{x-y+y-x-z+z}=\dfrac{dx+dy+dz}{0}"


"\Rightarrow dx+dy+dz=0"


Integrating and we get-


"x+y+z=c_1"


Also each ratio is equal to-


"\dfrac{dx-dy+dz}{x-y-y+x+z+z}=\dfrac{dx-dy+dz}{2(x-y+z)}"

Taking with 3 ratio of (1) we get


"\dfrac{dz}{z}=\dfrac{dx-dy+dz}{2(x-y+z)}"


"\Rightarrow \dfrac{dz}{z}=\dfrac{d(x-y+z)}{2(x-y+z)}"


"\Rightarrow 2\dfrac{dz}{z}=\dfrac{d(x-y+z)}{(x-y+z)}"


Integrating and we get-


"log(x-y+z)+logc_2=2logz"


"\Rightarrow logz(x-y+z)c_2=logz^2"


"\Rightarrow \dfrac{z^2}{x-y+z}=c_2"


The solution of the above equation is -


"\phi(c_1,c_2)=0"


"\phi(x+y+z,\dfrac{z^2}{x-y+z})=0"